# Designing AI-Resilient Assessments Using Interconnected Problems: A Theoretically Grounded and Empirically Validated Framework

### v1

@@ IV Results
IV-B Interconnectedness as a Differentiator: Open-Ended vs. Interconnected Projects
Our comparison of the Fall 2024 Predictive Analytics course (open-ended project) and the Fall 2025 offering (interconnected semi-open-ended project) reveals a more nuanced picture. While the average project scores were not dramatically different (91.47 vs. 89.66), the interconnected design led to a greater performance drop from the modular assignments and, critically, increased score variability. Figure 2 illustrates this comparison through box plots showing the distribution of scores.
@@
The performance drop from the modular average to the final project was larger in the interconnected course (3.07 points) than in the open-ended course (1.82 points). More importantly, the standard deviation of project scores was higher in the interconnected course (SD=21.93) compared to the open-ended course (SD=16.83), representing a 30% increase in absolute variability. This increased variance is not a flaw; it is evidence that the interconnected design created a more challenging and authentic assessment that could not be trivially solved by invoking an AI for each sub-problem. Table II provides a comprehensive overview of assessment performance across all four course offerings, highlighting the consistent patterns of performance gaps and correlations.
TABLE II: Summary of Assessment Performance Across Four Course Offerings
@@
Courses with Open-Ended Projects (Fall 2024)
@@
IV-C Predictive Validity of Assessments
We examined how well performance on modular assignments predicted performance on other assessment types. Table III presents a comprehensive correlation matrix for the Python course, revealing distinct patterns in how different assessment types relate to one another.
TABLE III: Correlation Matrix: Relationships Between Assessment Types in Python for Data Science Course (N=41)
Assessment Type Knowledge Homework Project Exam
Checks
Knowledge Checks 1.000
Homework 0.861*** 1.000
Project (Interconnected) 0.936*** 0.882*** 1.000
Proctored Exam 0.725*** 0.603*** 0.752*** 1.000
Modular Average 0.964*** 0.958*** 0.954*** 0.726***
*** $p<0.001$
Modular Average = (Knowledge Checks + Homework) / 2
The correlation between the modular average and exam scores was moderate (r = 0.726, p < .001), while the correlation between the modular average and the interconnected project scores was much stronger (r = 0.954, p < .001). This pattern is visualized in Figure 3, which presents scatter plots comparing these relationships. Panel (a) shows the tight clustering of points around the regression line for modular-project correlation, while panel (b) shows greater dispersion for modular-exam correlation.
@@ IV-C Predictive Validity of Assessments
Figure 3: Comparison of Predictive Validity: (a) Modular assessments strongly predict interconnected project performance (r=0.954), indicating both measure similar skills under different conditions. (b) Modular assessments moderately predict exam performance (r=0.726), suggesting exams may measure different skills or are less susceptible to AI inflation.
This finding suggests that interconnected projects, while more challenging, are a better measure of the same underlying skills being taught in the modular assignments. The strong correlation (r=0.954) indicates construct validity: both assessment types are measuring the same competencies, but the interconnected design does so under conditions that are less susceptible to AI inflation. In contrast, the moderate correlation with exams (r=0.726) may indicate that proctored exams are either testing a different set of skills (e.g., memorization, speed under pressure) or are simply less susceptible to AI influence due to their controlled environment. The visual contrast in Figure 3 makes this distinction immediately apparent, showing that interconnected projects preserve the pedagogical alignment with course content while adding AI-resilience.
@@ V Discussion
Our findings provide empirical support for the two theorems established in our theoretical framework. The 30-point performance gap between AI-permissive modular assignments and proctored exams (Table I) demonstrates what we term an AI inflation effect, where assessment scores may not accurately reflect student mastery when AI tools are available. While proctored exams are less susceptible to AI inflation, they may assess different competencies (e.g., time pressure, memorization) rather than the intended learning outcomes. Our data indicate that semi-structured, interconnected projects offer a third path: they preserve the authentic, workflow-based learning experiences that mirror industry practice while providing better differentiation of student ability than fully open-ended designs when AI tools are available.
More importantly, our results demonstrate that interconnectedness is a useful mechanism for creating AI-resilient assessments. Comparing two offerings of the same Predictive Analytics course—one with an open-ended project (Fall 2024, N=34) and one with an interconnected project (Fall 2025, N=22)—the interconnected design exhibited higher score variability (SD=21.93 vs. 16.83) despite similar mean scores (89.66 vs. 91.47, Table II). This increased variance is not a flaw; it is a feature. It indicates that the interconnected project was a better differentiator of student ability, creating a more challenging and authentic assessment that could not be trivially solved by invoking an AI for each sub-problem. This supports Theorem 1, which posits that forcing multi-step, stateful reasoning increases resilience to AI delegation.
Furthermore, the strong correlation between modular assessments and interconnected project scores (r=0.954, p<.001, Table III) suggests that both are measuring similar underlying skills. However, the interconnected design does so under conditions that are less susceptible to AI inflation, making it a more robust measure of those skills. In contrast, the moderate correlation between modular assessments and proctored exam scores (r=0.726, p<.001, Table III) suggests that exams may measure different competencies or are influenced by factors such as time pressure and test-taking ability. This addresses the core challenge of modern assessment: how to evaluate the skills taught in a curriculum in a way that is both authentic and AI-resilient.
Our findings also lend support to Theorem 2. The high scores and low variance on open-ended projects suggest that students (likely with AI assistance) may be choosing a path of least resistance, solving a simpler version of the problem than intended. The semi-open-ended, interconnected design, by contrast, constrains the problem space and forces engagement with the specific, intended challenges, resulting in a more reliable measure of competency.
@@ V-A A Design Procedure for Implementation
To translate these theoretical insights into actionable guidance for educators, we present a systematic design procedure for AI-resilient assessments (Algorithm 1). This procedure operationalizes Theorems 1 and 2 through six concrete steps that can be adapted to various computing education contexts.
@@
1:Step 1: Identify Core Workflow: $W\leftarrow$ Map $T$ to multi-stage workflow $\triangleright$ e.g., clean $\rightarrow$ EDA $\rightarrow$ model $\rightarrow$ evaluate
2:Step 2: Create Sequential Dependencies (Theorem 1): $stages\leftarrow$ Decompose $W$ where $input(s_{i})=output(s_{i-1})$
3:Step 3: Impose Semi-Structured Constraints (Theorem 2): Define deterministic success criteria $\triangleright$ e.g., F1 >0.90>0.90 on test set
4:Step 4: Embed Verifiable Checkpoints: For each stage, define intermediate output with validation $\triangleright$ e.g., zero nulls
5:Step 5: Introduce Domain Complexity: Embed challenges requiring contextual reasoning $\triangleright$ e.g., class imbalance
@@
7:return AA $\triangleright$ AI-resilient assessment design
The algorithm emphasizes several key principles. First, Step 2 directly operationalizes Theorem 1 by creating sequential dependencies that force multi-step reasoning and exceed typical AI context windows. Second, Step 3 implements Theorem 2 by constraining the solution space through semi-structured requirements, preventing AI from defaulting to easier problem variants. Third, Steps 4 and 5 introduce additional layers of complexity—verifiable checkpoints and domain-specific challenges—that further increase AI-resilience while maintaining pedagogical validity.
@@ VI Conclusion and Future Work
This paper makes three primary contributions to the field of computing education in the age of generative AI. First, we establish a theoretical foundation through two formal theorems that position interconnectedness and semi-structured design as key principles for creating AI-resilient assessments. Second, we validate this framework with empirical evidence from four university data science courses, demonstrating a substantial AI inflation effect in traditional modular assessments while showing that interconnected projects preserve construct validity and provide superior differentiation of student ability. Third, we operationalize these findings into a practical design procedure that translates theoretical principles into implementable assessment strategies.
Our work has immediate practical implications for educators. By shifting focus from the difficulty of individual tasks to the complexity of the connections between them, instructors can design assignments that preserve authentic, workflow-based learning experiences while naturally resisting trivial AI delegation. The framework we propose offers a systematic, theoretically-grounded alternative to ad-hoc solutions like banning AI or reverting to traditional in-person exams, which may assess different competencies than the intended learning outcomes.

### v3

@@ IV-A Descriptive Analysis: The AI Inflation Effect
The observed AI-driven grade inflation effect in modular take-home assessments, as discussed in Section IV-A, highlights the need for AI-resilient assessment design, as modular assessment materials no longer effectively gauge student learning progress.
IV-B Comparative Analysis
IV-B1 Interconnectedness as a Differentiator: Fall 2024–Fall 2025 Predictive Analytics Course Assessment Results Empirical Comparison
We compare the Fall 2024 and Fall 2025 offerings of the Predictive Analytics course to examine the effect of interconnected assessment design. Both courses shared identical syllabi and encouraged the use of AI tools. The Fall 2024 course project did not incorporate the interconnected design principles outlined in Proposition 1, whereas the Fall 2025 offering explicitly integrated interconnected assessment into the project design. Both offerings included regular homework assignments in addition to course projects. The interconnected design in Fall 2025 resulted in notably lower average project scores (M = 78.58) compared to Fall 2024 (M = 91.47). Critically, the interconnected design also increased score variability, with standard deviations of 30.42 (Fall 2025) versus 16.83 (Fall 2024), as shown in Figure 2. This reduction in average scores and increase in variance suggests that the interconnected assessment better differentiates student competency levels, whereas the open-ended design in Fall 2024 allowed more students to achieve uniformly high scores—potentially through AI assistance.
@@ IV-B Comparative Analysis
This increased variance is not a flaw; it is evidence that the interconnected design created a more challenging and authentic assessment that could not be trivially solved by invoking Generative AI tools for each sub-problem. Table II provides a comprehensive overview of assessment performance across all three course offerings, highlighting the consistent patterns of performance gaps and correlations.
TABLE II: Summary of Assessment Performance Across three Course Offerings
@@
Fall 2024
@@
IV-B2 Empirical Evidence Against Open-Ended Assessments in Computing Education
In the Fall 2024 offering of a predictive analytics course, the final project was structured as a fully open-ended task. Students were granted unrestricted freedom in topic selection, data sourcing, and analytical methodology, with the expectation that they would analyze U.S. Bureau of Labor Statistics (BLS) data spanning 2008–2023 [37]. The primary motivation for this design was to encourage independent exploration and creative application of statistical and machine learning techniques in an authentic, real-world setting.
In a subsequent 2025 offering of the course, the project was redesigned as a semi-open-ended assignment. This revised format introduced explicit constraints on topical scope, minimum required data volume, required modeling techniques, and a set of interconnected analytical questions intended to guide—rather than prescribe—students’ exploration.
Several consistent empirical patterns emerged from the fully open-ended Fall 2024 implementation. First, many student projects relied on datasets that were substantially smaller or less variable than anticipated, despite the availability of high-dimensional and longitudinal data from the BLS. Second, a majority of students independently converged on commonly used macroeconomic inflation datasets. This convergence suggests a preference for highly discoverable data sources with well-established analytical narratives, and it did not consistently yield the diversity or originality often attributed to fully open-ended assessment designs [27, 14, 25, 39].
By contrast, in the more constrained Fall 2025 semi-open-ended project, we observed substantially greater engagement with larger-scale datasets, a broader range of modeling techniques, and increased use of advanced methods, including natural language processing. Students demonstrated greater analytical depth while still exercising flexibility in model choice and interpretation.
Taken together, these observations provide empirical support for Proposition 2, which posits that, in computing education contexts, semi-open-ended project designs may be more effective than fully open-ended alternatives. Fully open-ended tasks appear to incentivize least-resistance solution paths—an effect that may be amplified under AI-assisted workflows—whereas semi-open-ended designs preserve student autonomy while reducing convergence toward trivial, overused, or narratively “safe” data choices.
IV-C Correlation Analysis
Proctored examinations in which generative AI use is explicitly prohibited are, by construction, AI-resilient. However, professional practice in industry routinely involves the use of generative AI tools to support problem solving, analysis, and implementation. As a result, proctored exams do not fully reflect authentic professional workflows, whereas project-based assessments that permit generative AI use more closely mirror real-world practice.
Motivated by this tension, we explicitly examine how student performance on interconnected, semi-open-ended project assessments correlates with performance on proctored exams. We find that performance on the interconnected projects exhibits a substantially stronger correlation with exam performance (Pearson r = 0.925) than does performance on modular knowledge checks (r = 0.671), as shown in Figure 3. This result suggests that the interconnected assessment design provides a more valid measure of student understanding.
@@ IV-C Correlation Analysis
Figure 3: Comparison of predictive validity: (a) Interconnected project scores are strongly correlated with proctored exam scores (r=0.925), indicating that both assessments measure similar underlying competencies under different conditions. (b) Modular assessments exhibit a more moderate correlation with exam performance (r=0.671), suggesting greater susceptibility to AI-related score inflation.
The difference between these correlation coefficients is practically meaningful. Project performance explains approximately 86% of the variance in exam scores ($R^{2}=0.86$), compared to only 45% ($R^{2}=0.45$) for the modular knowledge checks. This substantial gap indicates that interconnected assessments provide a more accurate representation of what students actually know and are able to do.
Taken together, these findings suggest that interconnected semi-open-ended project-based assessments—while more demanding for both students and instructors—serve as a stronger proxy for the same underlying competencies evaluated by proctored exams. The high correlation (r = 0.925) indicates strong construct alignment between the two assessment modalities, while the project-based format evaluates these competencies under conditions that more closely resemble authentic, AI-augmented professional practice.
@@ V Discussion
Our results provide convergent empirical evidence for the central claims of our framework: when generative AI tools are permitted, modular assessments are vulnerable to score inflation, and that interconnected, semi-open-ended designs offer a more AI-resilient alternative without reverting to fully proctored exams.
First, the large performance gap observed between AI-permissive modular assessments and proctored exams in the Python for Data Science and AI course demonstrates a pronounced AI inflation effect. Near-ceiling performance on modular assignments contrasts sharply with substantially lower exam scores, indicating that high modular scores no longer reliably reflect learning outcome when AI tools are available.
Second, comparisons across 2024 and 2025 Predictive Analytics course with identical syllabi offerings show that assessment interconnectedness materially alters score distributions under AI-permissive conditions. The interconnected project design yielded lower mean scores and substantially higher variance than the fully open-ended project, despite comparable course structure and encouragement of AI use. This increase in dispersion is a desirable property: it indicates stronger differentiation of student ability and reduced feasibility of decomposing the task into independently solvable subproblems delegated to AI. These findings directly support Proposition 1, which posits that multi-step, stateful dependency structures increase resistance to AI-mediated shortcutting. Furthermore, our empirical comparison of topical and methodological choices between the fully open-ended (2024) and semi-open-ended (2025) projects reveals no meaningful differences in creative direction, data modality, or analytical framing. However, the open-ended design was associated with narrower data coverage, convergence toward highly discoverable inflation datasets, and fewer modeling approaches per project. This pattern is consistent with Proposition 2, by introducing structural constraints while preserving analytical flexibility, the semi-open-ended interconnected design reduced this convergence, increased methodological depth, and produced more reliable signals of student competency.
Finally, correlation analyses reveal meaningful differences in alignment between assessment formats and fully AI-resilient proctored exams. Performance on the interconnected, semi-open-ended project exhibits a substantially stronger correlation with proctored exam outcomes than do modular assessments, indicating that the interconnected design captures similar underlying competencies while remaining compatible with AI-augmented workflows. In contrast, modular assessments—despite high average scores—show weaker alignment with exam performance, reinforcing the conclusion that they are more susceptible to AI-driven score inflation.
@@ V-A A Design Procedure for Implementation
To translate these insights into actionable guidance for educators, we present a systematic design procedure for AI-resilient assessments (Algorithm 1). This procedure operationalizes Propositions 1 and 2 through six concrete steps that can be adapted to various computing education contexts.
@@
1:Step 1: Identify Core Workflow: $W\leftarrow$ Map $T$ to multi-stage workflow
2:Step 2: Create Sequential Dependencies (Proposition 1): $stages\leftarrow$ Decompose $W$ where $input(s_{i})=output(s_{i-1})$
3:Step 3: Impose Semi-Structured Constraints (Proposition 2): Define deterministic success criteria
4:Step 4: Embed Verifiable Checkpoints: For each stage, define intermediate output with validation
5:Step 5: Introduce Domain Complexity: Embed challenges requiring contextual reasoning
@@
7:return AA
Algorithm 1 emphasizes several key principles. First, Step 2 directly operationalizes Proposition 1 by creating sequential dependencies that force multi-step reasoning and exceed typical AI context windows. Second, Step 3 implements Proposition 2 by constraining the solution space through semi-structured requirements, preventing AI from defaulting to easier problem variants. Third, Steps 4 and 5 introduce additional layers of complexity—verifiable checkpoints and domain-specific challenges—that further increase AI-resilience while maintaining pedagogical validity.
@@ VI Conclusion and Future Work
This paper makes three primary contributions to the field of computing education in the age of generative AI. First, we establish two formal propositions that position interconnectedness and semi-structured design as key principles for creating AI-resilient assessments. Second, we validate this framework with empirical evidence from three university data science courses, demonstrating a substantial AI inflation effect in traditional modular assessments while showing that interconnected projects preserve construct validity and provide superior differentiation of student ability. Third, we operationalize these findings into a practical design procedure that translates theoretical principles into implementable assessment strategies.
Our work has immediate practical implications for educators. By shifting focus from the difficulty of individual tasks to the complexity of the connections between them, instructors can design assignments that preserve authentic, workflow-based learning experiences while naturally resisting trivial AI delegation. The framework we propose offers a systematic alternative to ad-hoc solutions like banning AI or reverting to traditional in-person exams altogether.